# Assessing and improving reliability of neighbor embedding methods: a map-continuity perspective

### v1

@@ Abstract
Visualizing high-dimensional data is an important routine for understanding biomedical data and interpreting deep learning models. Neighbor embedding methods, such as t-SNE, UMAP, and LargeVis, among others, are a family of popular visualization methods which reduce high-dimensional data to two dimensions. However, recent studies suggest that these methods often produce visual artifacts, potentially leading to incorrect scientific conclusions. Recognizing that the current limitation stems from a lack of data-independent notions of embedding maps, we introduce a novel conceptual and computational framework, LOO-map, that learns the embedding maps based on a classical statistical idea known as the leave-one-out. LOO-map extends the embedding over a discrete set of input points to the entire input space, enabling a systematic assessment of map continuity, and thus the reliability of the visualizations. We find for many neighbor embedding methods, their embedding maps can be intrinsically discontinuous. The discontinuity induces two types of observed map distortion: “overconfidence-inducing discontinuity,” which exaggerates cluster separation, and “fracture-inducing discontinuity,” which creates spurious local structures. Building upon LOO-map, we propose two diagnostic point-wise scores—perturbation score and singularity score—to address these limitations. These scores can help identify unreliable embedding points, detect out-of-distribution data, and guide hyperparameter selection. Our approach is flexible and works as a wrapper around many neighbor embedding algorithms. We test our methods across multiple real-world datasets from computer vision and single-cell omics to demonstrate their effectiveness in enhancing the interpretability and accuracy of visualizations.
Keywords— Data visualization, t-SNE, UMAP, leave-one-out, perturbation, manifold learning, clustering, discontinuity
1 Introduction
@@
Data visualization plays a crucial role in modern data science, as it offers essential and intuitive insights into high-dimensional datasets by providing low-dimensional embeddings of the data. Classical methods for dimension reduction include principal component analysis (PCA) [1], multidimensional scaling [2], locally linear embedding [3], among others. For visualizing high-dimensional data, the last two decades have witnessed rising popularity of t-SNE [4] and UMAP [5], which are extensively used in, e.g., single-cell analysis [6, 7, 8] and feature interpretations for deep learning models [9, 10].
@@
The neighbor embedding methods [11, 12] are a family of visualization methods, which include t-SNE, UMAP, and LargeVis [13] as popular examples, that determine embedding points directly by solving a complicated optimization algorithm to minimize the discrepancy between similarities of input data points and those of the corresponding low-dimensional points. Given input data ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$, a neighbor embedding algorithm $\mathcal{A}$} computes the points $({\bm{y}}_{1},\ldots,{\bm{y}}_{n})={\mathcal{A}}({\bm{x}}_{1},\ldots,{\bm{x}}_{n})$ in the 2D plane, aiming to preserve the essential structures of ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$. Due to the algorithmic complexity, $\mathcal{A}$} is often used as a black-box visualization tool. This is very different from classical dimension reduction methods such as PCA, where a parametric mapping ${\bm{f}}_{{\bm{\theta}}}$ is determined and any input point $\bm{x}$} is embedded through ${\bm{y}}={\bm{f}}_{{\bm{\theta}}}({\bm{x}})$. For neighbor embedding methods, in contrast, there is no clearly defined embedding map for the “embedding points” ${\bm{y}}_{1},\ldots,{\bm{y}}_{n}$.
@@
Due to the lack of explicit embedding maps [14], neighbor embedding methods suffer from both conceptual and practical difficulties. For conceptual difficulties, there is no obvious sample-independent notion of embedding maps, since the embedding points ${\bm{y}}_{1},\ldots,{\bm{y}}_{n}$ depend on $n$ input points ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$ collectively, which makes it challenging to understand the correspondence between an input point ${\bm{x}}_{i}$ and an embedding point ${\bm{y}}_{i}$. Thus, it is unclear what structures the embedding points inherit from the input points, even in ideal settings where inputs are drawn from known distributions or simple manifolds.
For practical difficulties, it has been recently recognized that neighbor embedding methods often produce misleading results by creating severe distortion through the embedding maps and introducing spurious clusters in low-dimensional visualization [15, 16]. Moreover, neighbor embedding methods are sensitive to the choice of optimization algorithms [17], initialization schemes [18], and hyperparameters [18, 19], which may lead to inconsistent interpretations [20].
@@
Some progress has been made, but significant challenges remain. On the theoretical side, insights have been obtained regarding embedding stages [19, 21, 22], force-based interpretations [23], and visualization quality [21, 22, 24]. On the practical side, schemes of robust initialization and hyperparameter selection have been proposed [6, 16, 25, 26, 27]. To enhance the faithfulness of neighbor embedding methods, multiple diagnostic quantities have been proposed [6, 16, 25, 26]. In particular, Johnson et al. [28] and Xia et al. [27] propose algorithms to detect embedding points with severe neighborhood distortion, whereas Sun et al. [29] propose a dynamic approach to assess visualization uncertainty. However, due to a limited understanding of the origins and nature of the observed artifacts, most existing diagnostic methods offer only partial solutions and rely on ad hoc fixes, sometimes even introducing new artifacts (Figure 7).
@@
In this work, we propose a notion of embedding map—which we call LOO-map—induced by a given neighbor embedding method $\mathcal{A}$}. LOO-map is a mapping in the classical sense and approximates the properties of $\mathcal{A}$} around each embedding point. It is based on a classical strategy from statistics known as the leave-one-out (LOO) method, which posits that adding, deleting, or changing a single input point has negligible effects on the overall inferential results. Using LOO, we can decouple the pairwise interaction in the algorithm $\mathcal{A}$}: we add a new input point $\bm{x}$} to ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$ and freeze ${\bm{y}}_{1},\ldots,{\bm{y}}_{n}$ in the optimization problem, allowing only one free variable $\bm{y}$}. We call the resulting minimizer ${\bm{f}}({\bm{x}})$ the LOO-map, which satisfies the approximation $({\bm{y}}_{1},\ldots,{\bm{y}}_{n},{\bm{f}}({\bm{x}}))\approx{\mathcal{A}}({\bm{x}}_{1},\ldots,{\bm{x}}_{n},{\bm{x}})$. By design, the LOO-map $\bm{f}$} not only satisfies ${\bm{f}}({\bm{x}}_{i})\approx{\bm{y}}_{i}$ for all $i$’s, but also reveals the embedding point ${\bm{f}}({\bm{x}})$ of a potential new input point $\bm{x}$}. As such, LOO-map extends the mapping defined over the discrete input set $\{{\bm{x}}_{1},\ldots,{\bm{x}}_{n}\}$ to the entire input space.
@@
LOO-map provides a unified framework for both understanding recognized issues such as distance distortion [16, 30], low stability [29], and lack of neighborhood preservation [16, 31], and uncovering novel insights such as embedding discontinuity. In our view, discontinuities of ${\bm{f}}({\bm{x}})$ represent an extreme form of distortion that accompanies topological changes in the embedding space, e.g., connected clusters become separated and a uniform shape is broken into fractured pieces. In contrast, classical dimension reduction methods such as PCA do not suffer from map discontinuity since a continuous parametric map ${\bm{f}}_{\theta}({\bm{x}})$ is constructed explicitly. In this regard, embedding discontinuity is an innate issue of the family of neighbor embedding methods.
@@
Using LOO-map, we identify two types of observed distortion patterns, one affecting global properties of the embedding map and the other affecting local relationships. Both types of distortions are a consequence of discontinuities in ${\bm{f}}({\bm{x}})$ and can cause topological changes of the embedding structures.
@@
We propose two point-wise diagnostic scores, namely perturbation scores and singularity scores, to quantify the severity of the two types of map discontinuity at each embedding point. Our approach is flexible and works as a wrapper around many neighbor embedding algorithms (Supplementary file Section E.1) without requiring any label information. To illustrate the diagnostic scores, we provide two use cases: the first shows that perturbation scores detect out-of-distribution data (or data with distribution shift) in computer vision, and the second shows that singularity scores aid users in selecting hyperparameters in single-cell data analysis.
The proposed method is evaluated extensively on simulated datasets and multiple real-world datasets, including two benchmarking datasets from computer vision and four single-cell omics datasets (see Supplementary Table S1 for description). The R package implementing our proposed method is available at the GitHub repository:
https://github.com/zhexuandliu/NE-Reliability-MapContinuity.
2 Results
2.1 Overview of methods
We provide an overview of LOO-map and demonstrate the proposed two diagnostic scores (Figure 1).
@@ 2.1 Overview of methods
First, we introduce a general diagnosis strategy for discerning and analyzing discontinuity structure for neighbor embedding methods (e.g., t-SNE, UMAP). Given input data points ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$ in a potentially high-dimensional space, e.g., attribute vectors or feature vectors, we run an embedding algorithm such as t-SNE that returns embedding points ${\bm{y}}_{1},\ldots,{\bm{y}}_{n}$ in the 2D embedding space. Internally, a neighbor embedding method solves an optimization problem involving $O(n^{2})$ pairwise interaction terms between embedding points and yields embedding points $({\bm{y}}_{1},\ldots,{\bm{y}}_{n})={\mathcal{A}}({\bm{x}}_{1},\ldots,{\bm{x}}_{n})$. The LOO strategy rests on the assumption that none of the interaction terms is dominant so that perturbing any single input point has negligible effects overall on the embedding points. We verify this assumption extensively on both simulated datasets and real datasets (Table 1, Supplementary Table S2, Methods A.1). The LOO strategy allows us to add a new input $\bm{x}$} and optimize its corresponding $\bm{y}$} while freezing existing embedding points $({\bm{y}}_{j})_{j\leq n}$, and as a result, LOO-map is determined by solving a smaller optimization problem with only $O(n)$ effective terms due to frozen pair interaction between other embedding points. We identify the discontinuity points of ${\bm{f}}({\bm{x}})$ as the source of the observed distortions and artifacts.
@@
Then, we devise two diagnostic quantities (scores) that help to quantitatively assess the quality of embedding points. Both scores are computed for each embedding point without requiring any label information (Figure 1a). The first quantity, namely the perturbation score, quantifies how much an embedding point ${\bm{y}}_{i}$ moves when the input ${\bm{x}}_{i}$ is moderately perturbed, which helps to probe discontinuity of ${\bm{f}}({\bm{x}})$ from the input space. The second quantity, namely the singularity score, measures how sensitive an embedding point is to an infinitesimal input perturbation, thus providing insights into the ${\bm{f}}({\bm{x}})$ at each specific location ${\bm{x}}={\bm{x}}_{i}$. The two scores, as we will show below, are motivated by different considerations and reveal qualitatively distinct features of the visualizations (Figure 1b-d).
@@
Figure 1: Overview: assessment of embeddings by neighbor embedding methods with an example of image data. a We use a standard pre-trained convolutional neural network (CNN) to obtain features of image samples from the CIFAR10 dataset, and then visualize the features using a neighbor embedding method, specifically t-SNE. b Basic ideas of singularity scores and perturbation scores. c t-SNE tends to embed image features into separated clusters even for images with ambiguous semantic meanings (as quantified by higher entropy of predicted class probabilities by the CNN). Perturbation scores identify the embedding points which have ambiguous class membership but less visual uncertainty. d Incorrect choice of the hyperparameter leads to visual fractures (FI discontinuity), which is more severe with a smaller perplexity. We recommend choosing the perplexity no smaller than the elbow point.
@@
Finally, we demonstrate how our proposed scores can improve the reliability of neighbor embedding methods. Two additional use cases are provided in Sections 2.6 and 2.7. Consider the workflow in Figure 1a: we obtain high-dimensional features of image data from a deep learning model (e.g., ResNet-18 [32]) and use t-SNE to embed features into 2D points. We observe that a fraction of inputs with ambiguous (mixed) class membership are misleadingly embedded into well-separated clusters (Figure 1c), thereby creating overconfidence in the cluster structure. Indeed, both ground-truth class labels and label-informed entropy scores suggest that the visualization under-represents the uncertainty for mixed points, making them appear more distinct than they should be (Figure 1c, middle, right panels). Further examination of image examples confirms such an artifact of reduced uncertainty in the embedding space. As a diagnosis, we find that embedding points with high perturbation scores correlate well with such observed (OI) discontinuity.
@@
Our second diagnostic scores can help hyperparameter selection. A practical challenge of interpreting t-SNE embeddings is that results may be sensitive to tuning parameters. Indeed, we find that a small perplexity tends to induce small and spurious structures similar to fractures visually, which suggests the presence of local (FI) discontinuity in the LOO-map $\bm{f}$} (Figure 1d). Our singularity score captures such FI discontinuity as more high-scoring points emerge under a smaller perplexity. With this diagnosis, we recommend choosing a perplexity no smaller than the elbow point of the FI discontinuity curve.
2.2 Leave-one-out as a general diagnosis technique
@@
We start with a generic setup for neighbor embedding methods that encompasses SNE [33], t-SNE [4], UMAP [5], LargeVis [13], PaCMAP [15], among others. Let ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{n}]^{\top}\in\mathbb{R}^{n\times d}$ be the input data matrix, and ${\bm{Y}}=[{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}\in\mathbb{R}^{n\times p}$ be the matrix of embedding points we aim to determine, where $p$ can be 2 or 3. To find the embedding matrix $\bm{Y}$}, we solve an optimization problem
$\min_{{\bm{y}}_{1},\ldots,{\bm{y}}_{n}\in\mathbb{R}^{2}}\sum_{1\leq i<j\leq n}\mathcal{L}(w(\bm{y}_{i},\bm{y}_{j});v_{i,j}({\bm{X}}))+Z(\bm{Y})$ (1)
where $(v_{i,j})_{i<j}$ are the similarity scores calculated from $({\bm{x}}_{i})_{i\leq n}\in\mathbb{R}^{d}$ in the input space, and $w$ is a function of embedding points $({\bm{y}}_{i})_{i\leq n}$ in the 2D or 3D space. The similarity scores $(v_{i,j})_{i<j}$ are often calculated based on a Gaussian kernel, and $w$ takes the form of a heavy-tailed kernel. For example, in t-SNE algorithm (see Supplementary file Section E.2 for other neighbor embedding methods),
@@
$w_{i,j}=w(\bm{y}_{i},\bm{y}_{j})=(1+\|{\bm{y}}_{i}-{\bm{y}}_{j}\|^{2})^{-1},\qquad Z(\bm{Y})=\log\Big(\sum_{k\neq l}w(\bm{y}_{k},\bm{y}_{l})\Big).$ (2)
A fundamental challenge of assessing the embeddings is that we only know how discrete points—not the input space—are mapped, since the optimization problem is solved numerically by a complicated algorithm. Consequently, it is unclear if underlying structures (e.g., clusters, low-dimensional manifolds) in the input space are faithfully preserved in the embedding space.
Consider adding a new point $\bm{x}$} to existing data points. We may wish to fix ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$ and analyze how embedding points ${\mathcal{A}}({\bm{x}}_{1},\ldots,{\bm{x}}_{n},{\bm{x}})$ change as we vary $\bm{x}$}, thereby quantifying the mapping of $\bm{x}$} under $\mathcal{A}$}. As such, the embedding points would depend on all $n+1$ input points.
@@
LOO is a generic decoupling technique that allows us to isolate the changes of one embedding point versus the others [34, 35, 36, 37, 38, 39]. Rooted in the stability idea [40, 41, 42], LOO assumes that adding (or deleting/modifying) a single input point does not change embedding points significantly (Figure 2a). This assumption allows us to study the map ${\bm{x}}\mapsto{\mathcal{A}}({\bm{x}}_{1},\ldots,{\bm{x}}_{n},{\bm{x}})$ approximately. Consider the optimization problem in Eqn. 1 with $n+1$ inputs points ${\bm{x}}_{1},\ldots,{\bm{x}}_{n},{\bm{x}}$ where we treat $\bm{x}$} as the $(n+1)$-th point. Under the LOO assumption, when adding the new input point $\bm{x}$}, we can freeze the embedding matrix $\bm{Y}=[{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}$ and allow only one free variable $\bm{y}$} in the optimization problem. This simplification results in the LOO loss function defined by
@@
We define the LOO-map as ${\bm{f}}:{\bm{x}}\mapsto{\rm argmin}_{{\bm{y}}}L({\bm{y}};{\bm{x}})$ (Figure 2b). The LOO loss is motivated by the following observation: suppose $\begin{bmatrix}\widetilde{{\bm{Y}}}\\ \widetilde{\bm{y}}^{\top}\end{bmatrix}$ is the embedding of ${\bm{X}}_{+}=\begin{bmatrix}\bm{X}\\ \bm{x}^{\top}\end{bmatrix}$, i.e., it reaches the minimum of the original loss, then ${\bm{y}}=\widetilde{\bm{y}}$ is necessarily the minimizer of a partial loss involving the embedding point of the added point $\bm{x}$}:

### v2

@@ Abstract
Visualizing high-dimensional data is essential for understanding biomedical data and deep learning models. Neighbor embedding methods, such as t-SNE and UMAP, are widely used but can introduce misleading visual artifacts. We find that the manifold learning interpretations from many prior works are inaccurate and that the misuse stems from a lack of data-independent notions of embedding maps, which project high-dimensional data into a lower-dimensional space. Leveraging the leave-one-out principle, we introduce LOO-map, a framework that extends embedding maps beyond discrete points to the entire input space. We identify two forms of map discontinuity that distort visualizations: one exaggerates cluster separation and the other creates spurious local structures. As a remedy, we develop two types of point-wise diagnostic scores to detect unreliable embedding points and improve hyperparameter selection, which are validated on datasets from computer vision and single-cell omics.
11footnotetext: Department of Statistics, University of Wisconsin–Madison, Madison, WI, 53706, USA.22footnotetext: Department of Biostatistics, T.H. Chan School of Public Health, Harvard University, Boston, MA 02115, USA.33footnotetext: Department of Data Science, Dana-Farber Cancer Institute, Boston, MA 02115, USA.44footnotetext: Email: yiqiao.zhong@wisc.edu
Introduction
@@
Data visualization plays a crucial role in modern data science, as it offers essential and intuitive insights into high-dimensional datasets by providing low-dimensional embeddings of the data. For visualizing high-dimensional data, the last two decades have witnessed the rising popularity of t-SNE [1] and UMAP [2], which are extensively used in, e.g., single-cell analysis [3, 4, 5] and feature interpretations for deep learning models [6, 7].
@@
The neighbor embedding methods [8, 9] are a family of visualization methods, which include t-SNE, UMAP, and LargeVis [10] as popular examples, that determine embedding points directly by solving a complicated optimization algorithm to minimize the discrepancy between similarities of input points and those of the corresponding low-dimensional points. Given input data ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$, a neighbor embedding algorithm $\mathcal{A}$} computes the points $({\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n})={\mathcal{A}}({\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n})$ in the 2D plane, aiming to preserve the essential structures of ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$. Due to algorithmic complexity, $\mathcal{A}$} is often used as a black-box visualization tool.
[…]
@@
A key conceptual difficulty is the lack of sample-independent notion of embedding maps, since the embedding points ${\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}$ depend on $n$ input points ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$ collectively, which makes it challenging to understand the correspondence between an input point ${\mathbf{x}}_{i}$ and an embedding point ${\mathbf{y}}_{i}$. Thus, it is unclear what structures the embedding points inherit from the input points, even in ideal settings where inputs are drawn from known distributions or simple manifolds. The lack of continuous-space embedding maps leads to recent recognition that neighbor embedding methods often produce misleading results by creating severe distortion through the embedding maps and introduce spurious clusters in low-dimensional visualization [15, 16]. Moreover, neighbor embedding methods are sensitive to the choice of optimization algorithms [17], initialization schemes [18], and hyperparameters [18, 19], leading to inconsistent interpretations [20, 21].
@@
Some progress has been made to improve the reliability of these visualization methods, including insights on embedding stages [19, 22, 23], force-based interpretations [24], visualization quality [22, 23, 25], initialization schemes, and hyperparameter selection [3, 16, 26, 27, 28]. To enhance the faithfulness of neighbor embedding methods, multiple diagnostic approaches have been proposed [3, 16, 26, 27, 29, 30, 31]. However, most existing diagnostic methods offer only partial solutions and rely on ad hoc fixes, sometimes even introducing new artifacts.
In this work, we show that the manifold learning interpretation, which implicitly assumes a continuous mapping, is inaccurate. Our analyses reveal intrinsic discontinuity points in the embeddings that result in severe distortions. Our results imply that t-SNE and UMAP—which can induce topological changes to visualization—are fundamentally different from PCA and other parametric embedding methods.
@@
We address the conceptual difficulty by proposing a notion of embedding map—which we call LOO-map—induced by a given neighbor embedding method $\mathcal{A}$}. LOO-map is a mapping in the classical sense and approximates the properties of $\mathcal{A}$} around each embedding point. It is based on a well-established strategy from statistics known as the leave-one-out (LOO) method, which posits that adding, deleting, or changing a single input point has negligible effects on the overall inferential results. Using LOO, we can decouple the pairwise interaction in the algorithm $\mathcal{A}$}: we add a new input point $\mathbf{x}$} to ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$ and freeze ${\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}$ in the optimization problem, allowing only one free variable $\mathbf{y}$}. We call the resulting minimizer ${\mathbf{f}}({\mathbf{x}})$ the LOO-map, which satisfies the approximation $({\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n},{\mathbf{f}}({\mathbf{x}}))\approx{\mathcal{A}}({\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n},{\mathbf{x}})$. By design, the LOO-map $\mathbf{f}$} not only satisfies ${\mathbf{f}}({\mathbf{x}}_{i})\approx{\mathbf{y}}_{i}$ for all $i$’s, but also reveals the embedding point ${\mathbf{f}}({\mathbf{x}})$ of a potential new input point $\mathbf{x}$}. As such, LOO-map extends the mapping defined over the discrete input set $\{{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}\}$ to the entire input space.
@@
LOO-map offers a unified framework for understanding known issues like distance distortion [16, 32], low stability [30], and poor neighborhood preservation [16, 33], while also revealing new insights into embedding discontinuity. In our view, discontinuities of ${\mathbf{f}}({\mathbf{x}})$ represent an extreme form of distortion that accompanies topological changes in the embedding space, e.g., connected clusters become separated and a uniform shape is fractured into pieces. In contrast, classical dimension reduction methods such as PCA do not suffer from map discontinuity since a continuous parametric map ${\mathbf{f}}_{\theta}({\mathbf{x}})$ is constructed explicitly. In this regard, embedding discontinuity is an innate issue of the family of neighbor embedding methods.
@@
Using LOO-map, we identify two types of observed distortion patterns, one affecting global properties of the embedding map and the other affecting local relationships. Both types of distortion are a consequence of discontinuities in ${\mathbf{f}}({\mathbf{x}})$ and can cause topological changes in the embedding structures.
@@
We propose two types of point-wise diagnostic scores, namely perturbation scores and singularity scores, to quantify the severity of the two types of map discontinuity at each embedding point. Our approach is flexible and works as a wrapper around many neighbor embedding algorithms (Supplementary File Section A.1) without any label information.
In this work, we demonstrate the utility of our method through two use cases: detecting out-of-distribution data (or distribution shifts) in computer vision using the perturbation score, and selecting hyperparameters in single-cell data analysis using the singularity score. We evaluate our method on multiple simulated and real-world datasets (Supplementary Table S1, Methods). Comparisons with existing approaches show that our method achieves superior performance in detecting topological changes in embedding and hyperparameter selection. The R package implementing our method, along with a tutorial, is publicly available on GitHub:
https://github.com/zhexuandliu/MapContinuity-NE-Reliability.
Results
Overview of methods
We provide an overview of LOO-map and demonstrate the proposed two diagnostic scores (Fig. 1).
@@ Results
First, we introduce a general strategy to discern and analyze discontinuities in neighbor embedding methods (e.g., t-SNE, UMAP). Given input points ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$ in a potentially high-dimensional space, e.g., attribute vectors or feature vectors, an embedding algorithm $\mathcal{A}$} maps them to 2D points ${\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}$ by solving an optimization problem involving $O(n^{2})$ pairwise interaction terms. The LOO strategy assumes no dominant interaction term so that perturbing any single input point has negligible effects on the overall embedding. We extensively verify this assumption on simulated and real datasets (Table 1, Supplementary Table S2, Methods). By adding a new input $\mathbf{x}$} and optimizing its corresponding $\mathbf{y}$} while freezing $({\mathbf{y}}_{j})_{j\leq n}$, LOO-map reduces the optimization problem to only $O(n)$ effective interaction terms. We identify the discontinuity points of ${\mathbf{f}}({\mathbf{x}})$ as the source of the observed distortions and artifacts.
@@
Then, we devise two label-free point-wise diagnostic scores to quantitatively assess embedding quality (Fig. 1a). The first quantity, namely the perturbation score, quantifies how much an embedding point ${\mathbf{y}}_{i}$ moves when the input ${\mathbf{x}}_{i}$ is moderately perturbed, which helps to probe the discontinuity of ${\mathbf{f}}({\mathbf{x}})$ from the input space. The second quantity, namely the singularity score, measures how sensitive an embedding point is to an infinitesimal input perturbation, thus providing insights into ${\mathbf{f}}({\mathbf{x}})$ at each specific location ${\mathbf{x}}={\mathbf{x}}_{i}$. The two scores, as we will show below, are motivated by different considerations and reveal qualitatively distinct features of the visualizations (Fig. 1b-d).
@@
Figure 1: Overview: assessment of embeddings generated by neighbor embedding methods, illustrated with image data. a We use a standard pre-trained convolutional neural network (CNN) to obtain features of image samples from the CIFAR10 dataset, and then visualize the features using a neighbor embedding method, specifically t-SNE. b Basic ideas of singularity scores and perturbation scores. c t-SNE tends to embed image features into separated clusters even for images with ambiguous semantic meanings (as quantified by higher entropies of predicted class probabilities by the CNN). Perturbation scores identify the embedding points that have ambiguous class membership but less visual uncertainty. d An incorrect choice of perplexity leads to visual fractures (FI discontinuity), which is more severe with a smaller perplexity. We recommend choosing the perplexity no smaller than the elbow point.
@@
Finally, we demonstrate how our scores can improve the reliability of neighbor embedding methods. Following the workflow in Fig. 1a, we extract high-dimensional features of image data using a deep learning model (e.g., ResNet-18 [34]) and apply t-SNE for the 2D embedding. We observe that some inputs with ambiguous (mixed) class membership are misleadingly embedded into well-separated clusters (Fig. 1c), creating overconfidence in the cluster structure. Ground-truth labels and label-informed entropy scores confirm that the visualization under-represents the uncertainty for mixed points, making them appear more distinct than they should be (Fig. 1c). Further examination of image examples confirms such an artifact of reduced uncertainty in the embedding space. As a diagnosis, we find that embedding points with high perturbation scores correlate well with such observed (OI) discontinuity.
@@
Our second diagnostic score can help with hyperparameter selection. A practical challenge of interpreting t-SNE embeddings is that the results may be sensitive to tuning parameters. In fact, we find that a small perplexity tends to induce small spurious structures similar to fractures visually speaking, suggesting the presence of local (FI) discontinuity in the LOO-map $\mathbf{f}$} (Fig. 1d). Our singularity score captures such FI discontinuity as more high-scoring points emerge under smaller perplexities. With this diagnosis, we recommend choosing a perplexity no smaller than the elbow point of the FI discontinuity curve.
Leave-one-out as a general diagnosis technique
@@
We start with a generic setup for neighbor embedding methods that encompasses SNE [35], t-SNE [1], UMAP [2], LargeVis [10], PaCMAP [15], among others. First, we introduce basic mathematical concepts and their interpretations.
  • Input data matrix ${\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}]^{\top}\in\mathbb{R}^{n\times d}$: the input data to be visualized. Dimension $d$ may be large (e.g., thousands).
  • Embedding matrix ${\mathbf{Y}}=[{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}\in\mathbb{R}^{n\times p}$: the embedding points we aim to determine for visualization, where $p$ can be 2 or 3.
  • (Pairwise) similarity scores $(v_{i,j})_{i<j}$: a measure of how close two input points are in the input space, often calculated based on a Gaussian kernel.
  • (Pairwise) embedding similarity scores $(w_{i,j})_{i<j}$: a measure of how close two embedding points are, which takes the form of a heavy-tailed kernel (e.g., t-distribution). The computation often requires a normalization step.
  • (Pairwise) loss function $\mathcal{L}$: a measure of discrepancy between $v_{i,j}$ and $w_{i,j}$. An NE method minimizes this loss function over embedding points to preserve local neighborhood structures. The algorithms of NE methods aim to find the embedding $\mathbf{Y}$} by minimizing the total loss composed of the sum of the divergences between $v_{i,j}$ and $w_{i,j}$ of all pairs of points and a normalization factor $Z({\mathbf{Y}})$.
For convenience, We introduce a generic optimization problem that neighbor embedding methods aim to solve as follows:
$\min_{{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}\in\mathbb{R}^{2}}\sum_{1\leq i<j\leq n}\underbrace{\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}}_{j});v_{i,j}({\mathbf{X}}))}_{\text{unnormalized pairwise loss}}+\underbrace{Z({\mathbf{Y}})}_{\text{normalization factor}}.$ (1)
Particularly, in the t-SNE algorithm (see Supplementary Methods A.2 for other algorithms), we have,
@@
$w_{i,j}=w({\mathbf{y}}_{i},{\mathbf{y}}_{j})=(1+\|{\mathbf{y}}_{i}-{\mathbf{y}}_{j}\|^{2})^{-1},\quad Z({\mathbf{Y}})=\log\Big(\sum_{k\neq l}w({\mathbf{y}}_{k},{\mathbf{y}}_{l})\Big).$ (2)
A fundamental challenge of assessing the embeddings is that we only know how discrete points—not the input space—are mapped since the optimization problem is solved numerically by a complicated algorithm. Consequently, it is unclear if underlying structures (e.g., clusters, low-dimensional manifolds) in the input space are faithfully preserved in the embedding space.
Consider adding a new point $\mathbf{x}$} to existing data points. We may wish to fix ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$ and analyze how embedding points ${\mathcal{A}}({\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n},{\mathbf{x}})$ change as we vary $\mathbf{x}$}, thereby quantifying the mapping of $\mathbf{x}$} under $\mathcal{A}$}. However, the embedding points would depend on all $n+1$ input points, and require re-running the neighbor embedding algorithm for each new $\mathbf{x}$}.
@@
LOO is a generic decoupling technique that allows us to isolate the changes of one embedding point versus the others [36, 37, 38, 39, 40, 41]. We introduce the LOO assumption, the LOO loss function, and LOO-map as follows.
  • LOO assumption: adding (or deleting/modifying) a single input point does not change embedding points significantly (Figure 2a).
  • LOO loss function $L({\mathbf{y}};{\mathbf{x}})$: it consists of $n$ pairwise loss terms relevant to the newly added point $\mathbf{x}$}. We aim to determine the embedding $\mathbf{y}$} for $\mathbf{x}$} (Fig. 2b).
  • LOO-map $\mathbf{f}$}: it is defined as ${\mathbf{f}}:{\mathbf{x}}\mapsto{\rm argmin}_{{\mathbf{y}}}L({\mathbf{y}};{\mathbf{x}})$ for all possible input $\mathbf{x}$} (Fig. 2b). This definition allows us to examine the map property in the entire region.
Rooted in the stability idea [42, 43, 44], LOO assumes that adding (or deleting/modifying) a single input point does not change embedding points significantly (Fig. 2a). This assumption allows us to study the map ${\mathbf{x}}\mapsto{\mathcal{A}}({\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n},{\mathbf{x}})$ approximately. Consider the optimization problem in Equation 1 with $n+1$ inputs points ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n},{\mathbf{x}}$. Under the LOO assumption, when adding the new $(n+1)$-th input point $\mathbf{x}$}, we can freeze the embedding matrix ${\mathbf{Y}}=[{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}$ and allow only one free variable $\mathbf{y}$} in the optimization problem. More precisely, the mathematical formulation of LOO loss function is given by
@@
The LOO loss is motivated by the following observation: suppose $\begin{bmatrix}\widetilde{{\mathbf{Y}}}\\ \widetilde{\mathbf{y}}^{\top}\end{bmatrix}$ is the embedding of ${\mathbf{X}}_{+}=\begin{bmatrix}{\mathbf{X}}\\ {\mathbf{x}}^{\top}\end{bmatrix}$, i.e., it reaches the minimum of the original loss, then ${\mathbf{y}}=\widetilde{\mathbf{y}}$ is necessarily the minimizer of a partial loss involving the embedding point of the point $\mathbf{x}$}: